OpenStudy (anonymous):

A bag has 5 red marbles, 6 blue, and 4 black. What's the probabilty of a black, replacing it, then getting a blue? i REALLY NEED HELP

OpenStudy (anonymous):

choices are: 4/45 8/75 2/15 (this is the wrong answer) 1/9

OpenStudy (anonymous):

i am thinking it is A but i am not positive

OpenStudy (anonymous):

Okay thank you! I'm really not sure right now

OpenStudy (anonymous):

I have another one as well. John is choosing a password for his access to internet. He decided not to use 0 and M. Each letter or number may be used more than once. How many can he make with 2 letters and 4 digits?

OpenStudy (anonymous):

Choices are: 4,100,625 3,100,625(wrong answer) 3,100,500 4,200,625

OpenStudy (anonymous):

sorry i am stuck

OpenStudy (anonymous):

Me to girl! me to..

OpenStudy (anonymous):

This is Algebra 2.. I suck at it pretty much..

OpenStudy (agent0smith):

A bag has 5 red marbles, 6 blue, and 4 black. What's the probabilty of a black, replacing it, then getting a blue? First, find how many marbles there are total: 5+6+4 = 15 There's 4 black, so the chance of getting a black is 4/15 Since we're replacing the black, there's still 15 marbles when we pick a blue. There is 6 blues, so the chance of getting a blue is 6/15 To find the chance of getting a black then blue, multiply those together: \[\frac{ 4 }{ 15 } \times \frac{ 6 }{ 15 } = ?\]